OpenStudy (anonymous):

Verify the following identity: cot(x - pi/2)= -tanx

OpenStudy (anonymous):

lets consider a triangle  find tan a and cot b

OpenStudy (anonymous):

Thats easy...\[\cot (x-y)=cotx*coty+1/(coty-cotx)\]

OpenStudy (anonymous):

put y=pi/2

OpenStudy (anonymous):

we get,\[\cot(x-\pi/2)=cotx (cotpi/2)+1/(\cot \pi/2-cotx)\]

OpenStudy (anonymous):

this gives answer 1/-cotx=-tanx